1/6 of the balls in a box are red, 5/12 of them are blue and 2/5 of the remainder is black. If there are 10 black balls, how many more blue balls are there than red ones?


Answer by stanbon(75887)

1/6 of the balls in a box are red, 5/12 of them are blue and 2/5 of the remainder is black. If there are 10 black balls, how many more blue balls are there than red ones?
----------
red = 1/6
blue = 5/12
Remainder = 1-(2/12 + 5/12) = 5/12
-------
Equation:
(5/12)x = 10
x = 24 (total # of balls)
---------
# of red balls = (1/6)24 = 4
# of blue balls = (5/12)24 = 10
----
Ans:: 10-4 = 6
==============
Cheers,
Stan H.
---------------